A car is sitting at a stoplight, when the light turns green at time t-Q. The road is straight. The cars distance from the stoplight as a function of time is given by z (t) = 2.80 t2-0.1 10 ts where this formula gives # in meters for t in seconds. The next 5 questions are referring to this situation
D | Question 5 1 pts e) How much time elapses from when the car first started to move until the car comes to a stop? Give your answer in units of seconds, rounded to 4 significant figures
Answer #1

x(t) = 2.8*t^2 - 0.011*t^3

So,

v = d x(t) / dt = 5.6*t - 0.33*t^2

So, for the car to come to stop, v = 0

So,  5.6*t - 0.33*t^2 = 0

So, t = 16.97 s <--------- answer